OpenStudy (anonymous):

What is the slope of the line passing through the point (4, −3) and perpendicular to the line y = 2x + 1? Write your answer as a fraction. What is the slope of a line that passes through the point (−5, 3) and is parallel to a line that passes through (2, 13) and (−4, −11)?

OpenStudy (anonymous):

For a line with slope m, the slope of a line perpendicular to it is -1 / m.

OpenStudy (anonymous):

Can you just tell me both the answers please? I really dont understand :((

OpenStudy (anonymous):

Oh its alright I got the first one. but not the 2nd

OpenStudy (anonymous):

For the first question y = 2x + 1 is in the form of y = mx + c, which is called the slope-intercept form, where m is the slope. So, m = 2. The slope of perpendicular line is -1 / m which is -1 / 2 (answer).

OpenStudy (anonymous):

Parallel means the slopes are the same

OpenStudy (anonymous):

Using the formula m = (y2 - y1) / (x2 - x1) where (x1 , y1) and (x2 , y2) are 2 points on the line.

OpenStudy (anonymous):

So question 2 m = (-11 - 13) / (-4 - 2) = -24 / -6 = 4